OpenStudy (anonymous):

Identify each point as either (a Solution) to the system (or Not a solution) to the system of inequalities. (1, 12) (2, 10) (4, 5) (6, 8.25) (12, 8) (18, 6)

OpenStudy (texaschic101):

where are the inequalities ?

OpenStudy (anonymous):

I dont know, it just says Identify each point as either (a Solution) to the system (or Not a solution) to the system of inequalities.

OpenStudy (texaschic101):

if we do not know the system, then we cannot answer this question

OpenStudy (anonymous):

Here´s some more information but how they have it sectioned in my school I didn´t think it would apply. Tyler went to the supermarket to buy food for a food pantry. He has $36, and can carry up to 20 pounds of food in his backpack. Pasta costs $1 for a 1-pound package. Pasta sauce costs $3 for a 1.5 pound jar. Let x = the number of packages of pasta and y = the number of jars of pasta sauce.

OpenStudy (texaschic101):

x + y <= 20 (pasta + sauce is less then or equal to 20 lbs) x + 3y <= 36 ($1 pasta + $3 sauce = $36) I think these are your inequalities....are you supposed to use the answer choices that you posted in your original post ?

OpenStudy (texaschic101):

you might better get a second opinion on this

OpenStudy (anonymous):

yes, the choices I originally posted. All the answer choices you have write beside a solution or not a solution to the system of inequalities

OpenStudy (texaschic101):

I can go through your answer choices to see if they satisfy the inequalities...I am just not 100% sure that the inequalities are written correctly. That is what you need the second opinion on.

OpenStudy (texaschic101):

the first inequality is correct...I am just not sure about the second one

OpenStudy (anonymous):

that's how it's displayed on my work

OpenStudy (texaschic101):

how whats displayed on your work ? Are the inequalities displayed on your work ?

OpenStudy (anonymous):

These are displayed on my work. Its just asking which ones are a solution and the ones that are not a solution (1, 12) (2, 10) (4, 5) (6, 8.25) (12, 8) (18, 6)

OpenStudy (texaschic101):

I think I found my mistake... 1x + 1.5y <= 20 -- (1 lb pasta + 1.5 lb sauce is less then or equal to $20) 1x + 3y <= 36 -- ($1 pasta + $3 sauce is less then or equal to $36) I am not gonna do all of these, but I will show you how to do them. Each one of your answer choices are set up like this (x,y) now just sub your x's and y's into both inequalities and see if they satisfy them. They have to satisfy both of them for it to be a solution. I will do a couple for you... (1,12) 1x + 1.5y <= 20 1(1) + 1.5(12) <= 20 1 + 18 <= 20 19 <= 20 (correct) now check the other inequality.. 1x + 3y <= 36 1(1) + 3(12) <= 36 1 + 36 <= 36 37 <= 36 (incorrect) The first answer choice is NOT A SOLUTION because it did not satisfy BOTH inequalities (2,10) 1x + 1.5y <= 20 1(2) + 1.5(10) <= 20 2 + 15 <= 20 17 <= 20 (correct) now for the 2nd one 1x + 3y <= 36 1(2) + 3(10) <= 36 2 + 30 <= 36 32 <= 36 (correct) This one IS A SOLUTION because it satisfies both inequalities. It is not really very hard to do, just time consuming. Do you have any questions ?

OpenStudy (anonymous):

no thanks